# Giovanni Compiani Ilya Morozov Stephan Seiler

### v3

@@ 3 Validation with Experimental Data
We apply our method to data from a choice experiment to show it recovers substitution patterns more effectively than standard attribute-based methods. A key challenge is that we usually do not directly observe substitution patterns in the data, and thus do not have a “ground truth” to validate demand models. To address this, we design and implement a choice experiment that measures second-choice diversion ratios, which directly reflect substitution.
@@ 3.1 Experiment Design
Each participant completes two choice tasks. In the first choice task, a participant chooses one book from a set of ten alternatives. We instruct participants to choose books they would purchase if faced with this selection in a real bookstore. We limit the number of options to ten to ensure participants pay attention to most options, which allows us to abstract away from limited consideration. After a participant selects a book from the list of ten options, we remove this book from the list. The participant then proceeds to a second choice task, where they choose among the remaining nine books.
Figure 1 shows the choice task as presented to participants. To give all models a fair chance at capturing substitution, we display attributes (author, year, genre, pages), cover images, and texts (plot description and five reviews), collected from Amazon product pages. We take all books from Amazon’s bestseller lists, sampling them in a way that does not bias our model comparisons in favor of any particular specification (see Appendix B for details). Lastly, we randomize books’ rankings and prices across participants, keeping them fixed across choice tasks for the same participant. [^14] Prices were drawn from a discrete uniform distribution ranging from \$3 to \$7. This double randomization generates clean variation for estimating substitution patterns (Berry and Haile, 2024). [^15] Although all participants technically see the same set of ten books in the first choice task, they rarely choose books that are displayed near the bottom of the list. Therefore, by randomizing the order in which books are displayed to participants, we vary their effective choice sets.
The book category is well-suited for our analysis because participants are likely to consider both structured attributes (e.g., genre) and unstructured information (e.g., plot descriptions). It is also unclear a priori whether substitution patterns are better predicted by images or texts—while texts describe rich plot details, cover images contain clear cues of genres, and whether the book is fiction or non-fiction (see Section 3.3 for a detailed discussion).
@@ 3.2 Model Specifications and Estimation Results
We compare our approach against two benchmark models: the plain logit model and a standard logit model with random coefficients model on attributes. All three models fall into a framework where the indirect utility of participant $i$ from book $j$ is
@@
Here $x_{j}$ are observed book attributes: genre dummies, publication year, and length in pages—all attributes participants observe in the choice tasks; $PC_{j}$ are principal components extracted from embeddings, $\text{price}_{ij}$ and $\text{rank}_{ij}$ are the price and position of book $j$ in participant $i$’s choice task, $\delta_{j}$ are product fixed effects, and $\varepsilon_{ij}$ are i.i.d. taste shocks following a Type I Extreme Value distribution. We include $\text{rank}_{ij}$ in the model because our experiment induces random variation in the placement of books. [^17] We do not include a random coefficient on $\text{rank}_{ij}$ in any model specification because our focus is on estimating substitution patterns using observed characteristics or principal components. This variation gives us an additional exogenous shifter of choices, which helps us identify substitution patterns in a manner similar to how price variation does.
@@
    Plain Logit. We include neither attributes $x_{j}$ nor principal components $PC_{j}$, estimating a constant price coefficient $\alpha_{i}=\alpha$, rank coefficient $\gamma$ and fixed effects $\delta_{j}$.
Table 1 compares in-sample $AIC$. The best mixed logit with attributes includes random coefficients on the number of pages and year, reducing $AIC$ by $16.0$ relative to plain logit. The best model with unstructured data, Reviews USE, puts random coefficients on the first two principal components. This model fits better than any other considered specification, reducing $AIC$ by $24.8$ relative to plain logit and by $8.8$ relative to the best attribute-based logit. While this superior fit is reassuring, it does not guarantee better counterfactual performance. We therefore turn to predicting second choices counterfactually.
@@ 3.3 Validation Using Second-Choice Data
To evaluate demand models, we let them confront a difficult counterfactual prediction problem—predicting second choices, i.e. which books participants switch to when their most preferred option becomes unavailable. A model can only predict second choices well if it has accurately captured the true substitution patterns. Predicting second choices is also directly relevant in antitrust where second-choice diversion ratios are often used to measure substitutability and the intensity of price competition (Conlon and Mortimer, 2021).
Operationally, we first estimate each model via Maximum Likelihood Estimation (MLE) using first choices only. We then use the estimated model to predict counterfactual second-choice diversion ratios, $\hat{s}_{j\rightarrow k}$, defined as the probability that the participant chooses product $k$ in the second choice task conditional on having chosen book $j$ in the first choice task. We compare these predictions with diversions observed in the data, $s_{j\rightarrow k}$, computing $RMSE$ as:
@@
Since we do not use second-choice data in estimation, lower $RMSE$ indicates that the model performs better at counterfactual predictions of second choices. Appendix D provides further details on how we compute diversions $s_{j\rightarrow k}$ and $\hat{s}_{j\rightarrow k}$.
Figure 2: Comparison of models in terms of counterfactual RMSERMSE on second choices. The two benchmarks in the first panel are the plain logit without random coefficients and the best-fitting mixed logit with random coefficients on observed attributes. The remaining specifications correspond to mixed logit models with random coefficients on principal components extracted from image or text embeddings.
@@
All four image models outperform the plain logit model, with the lowest-$AIC$ model, InceptionV3, reducing $RMSE$ by 7.0% relative to plain logit. To understand why images predict substitution, consider the book covers used in our study (see Figure 3). Within the same genre, book covers often share similar design elements. For example, the covers of all three fantasy books use dark, muted color palettes with metallic accents, include symbolic elements such as skulls and swords that convey danger or peril, and feature natural objects like twisted vines and golden roses. Similarly, the self-help books have minimalistic layouts and consistent color schemes, such as black-and-white text on yellow backgrounds. Thus, image embeddings partly encode books’ genres, which correlate with substitution patterns.
@@
First, performance tends to improve when we move from simple bag-of-words models to the more advanced USE and ST models. This is particularly true for descriptions and reviews, which contain detailed information about book plots but do not always describe them in the same words or phrases. Therefore, extracting substitution patterns from text requires natural language models that can accurately measure semantic similarity.
Second, performance improves as we include richer text data. For instance, the USE model reduces $RMSE$ by 4.3% with book titles, 17% with descriptions, and 23% with reviews relative to the plain logit. This makes intuitive sense. Although book titles are brief, they often contain subtle genre cues. For example, mystery titles frequently hint at secrets (Please Tell Me) or characters in confined situations (Housemaid, Inmate), whereas fantasy titles often include words that evoke a gothic atmosphere (Serpent, Wings, Ravens, Ruin). Descriptions provide even clearer genre signals, emphasizing unexpected twists such as hidden identities, past relationships, and betrayals. Reviews go even further: consumers summarize plots while also expressing opinions. When reviewing mystery novels, for instance, some readers praise cliffhangers and intriguing twists while others critique pacing issues like slow starts or rushed endings. Thus, it is unsurprising that reviews provide more information about substitution than descriptions, which, in turn, are more informative than titles.
@@
If text, images, and attributes provides distinct signals of substitution, combining them might improve fit and counterfactual performance. To explore this, we start from the selected Review USE model and attempt to extend it in two ways: (1) by adding a combination of observed product attributes with random coefficients, or (2) by adding a combination of principal components from InceptionV3, the lowest-$AIC$ image model. Neither extension improves $AIC$ or $RMSE$. This result suggests information across data types is highly correlated, with text providing the strongest signal of substitution.
@@
To better understand the variation captured by embeddings, in Figure 4 we show book locations in the space of the two principal components selected into our best-fitting Review USE model. These principal components align with intuitive substitution patterns: the first one (horizontal axis) separates non-fiction on the left from fiction on the right, while the second (vertical axis) further distinguishes science fiction from mystery. Crucially, the variation in these principal components goes beyond separating genres. For example, The Housemaid and The Inmate are by the same author, and The Serpent & The Wings and The Ashes & The Star Cursed King are from the same book series. The two principal components detect this similarity from consumers’ reviews of these books. As evident from $RMSE$ comparisons in Figure 2, this additional variation helps recover substitution patterns better and improves counterfactual predictions.
@@
Although we select models based on first-choice $AIC$, our experimental data allows us to verify whether this selection algorithm indeed chooses specifications with the best counterfactual performance. Two findings support our choice of $AIC$ as a model selection criterion. First, across all specifications with principal components considered by our model selection algorithm, the correlation between first-choice $AIC$ and counterfactual second-choice $RMSE$ is 0.78. Further, $AIC$ selects the specification that has the lowest second-choice $RMSE$ across all considered specifications. Using $BIC$ instead of $AIC$ selects the same specification, as does a five-fold cross-validation procedure on the first choices. This reassures us that the key results are robust to the choice of model selection method.
3.4 Implications for Pricing: Merger Simulations
To illustrate how estimated substitution patterns can influence counterfactuals of interest, we conduct simulations of horizontal mergers—a natural application of our approach. Antitrust agencies routinely use demand models to assess whether a hypothetical merger would lead to a significant price increase (Federal Trade Commission, 2022). If the merging firms’ products are close substitutes, the merger creates strong upward pricing pressure, as the firm can “recapture” some consumers after raising prices. Therefore, predicting how a merged firm would set prices requires accurate estimates of substitution patterns.
Figure 5: Results of hypothetical merger simulations. For each simulated merger of Dopamine Detox with one other book (horizontal axis), the figure shows the average price increase among the two merging books. The horizontal dashed line represents a hypothetical policy where the decision-maker challenges all mergers expected to raise prices by at least 5%. Appendix Table A2 reports the exact price increase estimates used to construct this graph.
For each pair of books, we compute their prices under two scenarios: (a) when the books are owned by separate publishers competing in a Bertrand-Nash equilibrium, and (b) when both books are owned by the same publisher setting prices by solving a joint first-order condition. In both cases, publishers take the prices of the other eight books as given, fixed at \$5, and have zero marginal costs. [^19] We fix the prices of the other books at \$5, the average value in the experimental dataset. We do not optimize prices of other books because the model does not include an outside option. Consumer choices remain unchanged if all prices increase by the same amount, thus leading to multiple equilibria.
We recognize that limiting the choice set to ten books makes our analysis somewhat artificial, as major publishers typically manage vast assortments with hundreds of thousands of titles. Nevertheless, our simulations offer a natural way to evaluate how estimated substitution patterns translate into pricing decisions. This application is also policy-relevant given the high-profile mergers among major book publishers in the past few decades, which have drawn attention and regulatory scrutiny. [^20] In 2022, a federal judge blocked the proposed merger between Penguin Random House and Simon & Schuster, citing concerns that it would stifle competition (U.S. Department of Justice, 2022). This merger would have reduced the “Big Five” publishers to the “Big Four” and significantly increased the concentration of the publishing market. Prior to that, Penguin Books and Random House merged in 2013 to form Penguin Random House—the largest trade book publisher globally.
Figure 5 illustrates the predicted price increase resulting from the joint ownership of Dopamine Detox with each of the other books. We select Dopamine Detox as an example where our approach outperforms alternative models in capturing substitution patterns (see Table 2). For each simulated merger, Figure 5 reports the average price increase across the two merging books.
As discussed above, the plain logit model fails to capture strong within-genre substitution. Consequently, it incorrectly identifies as the closest substitutes for Dopamine Detox the three mystery books, rather than other self-help ones. This misclassification affects the predicted price changes in merger simulations. Specifically, the plain logit model overstates the price increase when Dopamine Detox is merged with the mystery book Please Tell Me or The Housemaid, while understating the price increase when merged with its true closest substitutes, The Art of Letting Go or Don’t Believe Everything You Think. These discrepancies are substantial: for within-genre mergers, the plain logit predicts modest price increases of 3% and 3.7%, whereas our review-based model estimates them to be 5.6% and 8.9%—approximately 2 to 2.5 times higher (see Appendix Table A2).
@@ 4 Application to Online Retail Data
Next, we apply the same approach to choice data from several online markets. Our first goal is to show that this approach can be applied broadly across categories without being tailored. Our second goal is to identify which text and image data best predict substitution patterns in various categories, thus offering practical guidance on what types of unstructured data researchers should collect for demand estimation.
@@ 4.1 Data
We combine purchase data with image and text data we collected from Amazon product detail pages (Figure 6). We use the default product photo for image embeddings. For text embeddings, we use product titles, product descriptions (i.e., bullet points describing the product), and 100 most recent reviews for each product. In estimation, we treat text and image embeddings as fixed over time. While sellers could theoretically modify images and textual descriptions to reposition their products in response to unobserved demand shocks, we find that product images, titles, and descriptions very rarely change over time (see Appendix F).
We apply our method to 40 product categories spanning clothing (shirts, blouses, underwear, sleepwear), household goods (paper towels, trash bags, batteries), office supplies (pens, markers, printing paper), groceries (tea, coffee, bottled water), pet food (wet food, treats), electronics (tablets, monitors, headphones, memory cards, media players), and video games (PC, Nintendo, Xbox, PlayStation). We select the 15 most-purchased products in each category and retain only those categories where these products collectively account for at least 2,000 purchases. This criterion ensures that we observe enough purchases to estimate both product-fixed effects and substitution patterns. [^22] We omit products purchased fewer than 10 times and those with missing price data, which means some categories in our sample include fewer than 15 products. The average product price in selected categories is \$43.
For four electronics categories, we collect standard attributes from product detail pages to compare our approach with attribute-based mixed logit. Specifically, we collect 18 attributes for “Tablets,” 13 attributes for “Monitors,” 20 attributes for “Memory Cards,” and 14 attributes for “Headphones” (see Appendix E for a full list).
@@ 4.2 Estimation Results
We apply our approach to each product category and compare it to the plain logit model. In the four electronics categories for which we collected attribute data, we also compare it to the attribute-based mixed logit model. Since we have a relatively large number of attributes in these categories, we reduce their dimensionality via PCA and apply Algorithm 1 for model selection. [^23] We normalize each attribute to have mean zero and variance one before applying PCA.
@@
Appendix Table A4 shows results from the electronics categories with attribute data. In all four categories, unstructured data meaningfully improves model fit relative to observed attributes, suggesting that our approach captures information about substitution patterns beyond that contained in standard attributes. This finding is particularly noteworthy given that one might expect technical specifications of electronics products—captured by our extensive list of attributes—to be highly relevant for consumer choices.
Next, we examine the estimated substitution patterns. While we do not observe the “ground truth” diversion ratios as we did in the experiment, we can still assess how estimated diversion ratios deviate from those in the plain logit model. A well-known limitation of plain logit is that diversion ratios depend only on market shares and not on attribute similarities, so this model often fails to identify close substitutes, producing overly flat diversions (Conlon et al., 2023). If our approach recovers substitution better, it should produce more variable diversion ratios.
Figure 7 plots the estimated diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$. Consistent with the intuition above, the plain logit yields relatively small diversion ratios, about 22% on average, while our approach increases this average substantially to 47%. In some categories, our approach estimates diversion ratios as high as 60-80%, which the plain logit does not produce. These results confirm that our method generates significantly more variable diversion ratios, suggesting it better identifies which products are close substitutes.
Apart from showing greater variability, the diversion ratios predicted by our method are also more intuitive. For example, Appendix Tables A5 and A6 report predicted diversions for tablets. Because the plain logit relies on product fixed effects, it predicts that if any tablet is removed, consumers substitute to the most popular alternatives (e.g., Fire 7 or Fire HD10), producing unrealistically flat diversions. By contrast, the Description ST model chosen by our model selection algorithm yields more intuitive diversions: the two kids’ tablets are estimated to be close substitutes (Fire Kids and Fire HD8 Kids), the two iPads are close substitutes (iPad 9.7 and iPad 10.2), and so on.

### v4

@@ 3 Validation with Experimental Data
Next, we validate our approach empirically. A key challenge is that substitution patterns are unobserved, so there is typically no ground truth against which to test model predictions. To address this challenge, we design a choice experiment that provides such a ground-truth measure of substitution.
@@ 3.1 Experiment Design
Each participant completes two choice tasks. In the first choice task, a participant chooses one book from a set of ten alternatives. We instruct participants to choose books they would purchase if faced with this selection in a real bookstore. We limit the number of options to ten to ensure participants pay attention to most options, which allows us to abstract away from limited consideration. After a participant selects a book, we remove it from the list. The participant then proceeds to a second choice task, where they choose among the remaining nine books.
Figure 1 shows the choice task as presented to participants. To give all models a fair chance at capturing substitution, we display attributes (author, year, genre, pages), cover images, and texts (plot description and five reviews), collected from Amazon product pages. [^10] We collect the first five reviews displayed on each product detail page at the time of data collection. We take all books from Amazon’s bestseller lists, sampling them in a way that does not bias our model comparisons in favor of any particular specification (see Appendix B for details). Lastly, we randomize book rankings and prices across participants, keeping them fixed across choice tasks for the same participant. [^11] Prices were drawn from a discrete uniform distribution ranging from \$3 to \$7. This variation allows us to cleanly estimate substitution patterns from only first choices and reserve second-choice data for validation (Berry and Haile, 2024).
The book category is well-suited for our analysis because participants are likely to consider both structured attributes (e.g., genre) and unstructured information (e.g., plot descriptions). It is also unclear a priori whether substitution patterns are better predicted by images or texts—while texts describe rich plot details, cover images contain clear cues of genres, and whether the book is fiction or non-fiction (see Section 3.3 for further discussion).
@@ 3.2 Model Specifications and Estimation Results
We compare our approach against two benchmark models: the plain logit model and a mixed logit model with attributes. All three models fall into a framework where the indirect utility of participant $i$ from book $j$ is
@@
Here $x_{j}$ are observed book attributes: genre dummies, publication year, and length in pages—all attributes participants observe in the choice tasks; $PC_{j}$ are principal components extracted from embeddings, $\text{price}_{ij}$ and $\text{rank}_{ij}$ are the price and position of book $j$ in participant $i$’s choice task, $\delta_{j}$ are product fixed effects, and $\varepsilon_{ij}$ are i.i.d. taste shocks following a Type I Extreme Value distribution. We include $\text{rank}_{ij}$ in the model because our experiment induces random variation in the placement of books. [^13] We do not include a random coefficient on $\text{rank}_{ij}$ in any model specification because our focus is on estimating substitution patterns using observed characteristics or principal components. This variation gives us an additional exogenous shifter that helps us identify substitution patterns, analogous to price variation.
@@
    Plain Logit. We include neither attributes $x_{j}$ nor principal components $PC_{j}$, estimating only a constant price coefficient $\alpha_{i}=\alpha$, rank coefficient $\gamma$, and fixed effects $\delta_{j}$.
Table 1 compares in-sample $AIC$. The best mixed logit specification with attributes includes random coefficients on the number of pages and year, reducing $AIC$ by $16.0$ relative to plain logit. The best model with unstructured data, Reviews USE, puts random coefficients on the first two principal components. This model fits better than any other considered specification, reducing $AIC$ by $24.8$ relative to plain logit and by $8.8$ relative to the best attribute-based logit. While this superior fit is reassuring, it does not guarantee better counterfactual performance. We therefore turn to predicting second choices counterfactually.
@@ 3.3 Validation Using Second-Choice Data
We assess how well different demand models predict second choices counterfactually. Because second choices reveal which books participants choose when their preferred option is unavailable, predicting them accurately requires a model that captures substitution patterns well. Further, predicting second choices is directly relevant in antitrust where second-choice diversion ratios are often used as measures of substitutability and price competition (Conlon and Mortimer, 2021).
We first estimate each model via Maximum Likelihood Estimation (MLE) using only first choices. We then use the estimated model to predict counterfactual second-choice diversion ratios, $\hat{s}_{j\rightarrow k}$, defined as the probability that the participant chooses product $k$ in the second choice task conditional on having chosen book $j$ in the first choice task. We compare these predictions with diversions observed in the data, $s_{j\rightarrow k}$, computing $RMSE$ as:
@@
Since we do not use second-choice data in estimation, lower $RMSE$ indicates better performance at counterfactual predictions. Appendix D provides further details on how we compute diversions $s_{j\rightarrow k}$ and $\hat{s}_{j\rightarrow k}$.
Figure 2: Comparison of models in terms of counterfactual RMSERMSE on second choices. The two benchmarks in the first panel are the plain logit without random coefficients and the lowest-$AIC$ mixed logit with random coefficients on observed attributes. The remaining specifications correspond to mixed logit models with random coefficients on principal components extracted from image or text embeddings.
@@
All four image models outperform the plain logit model, with the selected (i.e., lowest-$AIC$) model, InceptionV3, reducing $RMSE$ by 7.0% relative to plain logit. To understand why images predict substitution, consider the book covers used in our study (see Figure 3). Within the same genre, book covers often share similar design elements. For example, the covers of all three fantasy books use dark, muted color palettes with metallic accents, include symbolic elements such as skulls and swords that convey danger or peril, and feature natural objects like twisted vines and golden roses. Similarly, the self-help books have minimalistic layouts and consistent color schemes, such as black-and-white text on yellow backgrounds. Thus, image embeddings partly encode books’ genres, which correlate with substitution patterns.
@@
First, performance improves as we move from simple bag-of-words models to the more advanced USE and ST models. This is particularly true for descriptions and reviews, which contain detailed information about book plots but do not always use the same words or phrases. Therefore, extracting substitution patterns from text requires natural language models that can accurately measure semantic similarity.
Second, performance improves as we include richer text data. For instance, relative to the plain logit, the USE model reduces $RMSE$ by 4.3% using titles, 17% using descriptions, and 23% using reviews. This makes intuitive sense. Take mystery books, for example. Titles may contain subtle genre cues, hinting at characters in confined situations (Housemaid, Inmate). Descriptions reveal further plot details, such as hidden identities, past relationships, and betrayals. Lastly, customer reviews contain the richest information about a book’s style and pacing: for instance, readers may praise cliffhangers and intriguing twists, or critique pacing issues like slow starts or rushed endings.
@@
If text, images, and attributes provide distinct signals of substitution, combining them might improve fit and counterfactual performance. To explore this, we start from the selected Review USE model and attempt to extend it in two ways: (1) by adding a combination of observed product attributes with random coefficients, or (2) by adding a combination of principal components from InceptionV3, the lowest-$AIC$ image model. In each case, we evaluate all subsets of added variables and estimate random coefficients for these added variables, in addition to those already included in the selected Review USE model. [^16] We chose this approach because including all principal components from unstructured data and all observed attributes in the candidate set for Algorithm 1 would be too computationally burdensome. We find that in all evaluated specifications, the variances of the added random coefficients are zero. As a result, $AIC$ increases and second-choice $RMSE$ does not improve. This result suggests information is highly correlated across data types, with text providing the strongest signal of substitution.
@@
To better understand the variation captured by embeddings, in Figure 4 we show book locations in the space of the two principal components selected into our best-performing Review USE model. These principal components align with intuitive substitution patterns: the first one (horizontal axis) separates non-fiction on the left from fiction on the right, while the second (vertical axis) further distinguishes science fiction from mystery. Crucially, the variation in these principal components goes beyond separating genres. For example, The Housemaid and The Inmate are by the same author, and The Serpent & The Wings and The Ashes & The Star Cursed King are from the same book series. The two principal components detect this similarity from consumers’ reviews of these books. As evident from $RMSE$ comparisons in Figure 2, this additional variation helps recover substitution patterns better and improves counterfactual predictions.
@@
Finally, to illustrate how estimated substitution patterns affect counterfactuals, we perform a stylized merger simulation in Appendix E. We find that the selected review-based model predicts substantially different post-merger prices than benchmark models.
@@
Although we select models based on first-choice $AIC$, our experimental data allows us to verify whether this selection algorithm indeed chooses specifications with the best counterfactual performance. Two findings support our choice of $AIC$ as a model selection criterion. First, across all specifications with principal components considered by our model selection algorithm, the correlation between first-choice $AIC$ and counterfactual second-choice $RMSE$ is 0.78. Second, $AIC$ selects the specification that has the lowest second-choice $RMSE$ across all considered specifications.
Lastly, using $BIC$ instead of $AIC$ selects the same specification, as does a five-fold cross-validation procedure on the first choices. This reassures us that the key results are robust to the choice of model selection method.
@@ 4 Application to Online Retail Data
Next, we apply our approach to choice data from several online markets. Our first goal is to show that this approach can be applied broadly across categories without being tailored. Our second goal is to determine which types of unstructured data best predict substitution patterns in various categories and to offer practical guidance on what data researchers should collect for demand estimation.
@@ 4.1 Data
We combine purchase data with image and text data we collected from Amazon product detail pages (Figure 5). We use the default product photo for image embeddings. For text embeddings, we use product titles, product descriptions (i.e., bullet points describing the product), and 100 most recent reviews for each product. In estimation, we treat text and image embeddings as fixed over time. While sellers could theoretically modify images and textual descriptions to reposition their products, we find that product images, titles, and descriptions very rarely change over time (see Appendix G).
We apply our method to 40 product categories spanning clothing (shirts, blouses, underwear, sleepwear), household goods (paper towels, trash bags, batteries), office supplies (pens, markers, printing paper), groceries (tea, coffee, bottled water), pet food (wet food, treats), electronics (tablets, monitors, headphones, memory cards, media players), and video games (PC, Nintendo, Xbox, PlayStation). We select the 15 most-purchased products in each category and retain only those categories where these products collectively account for at least 2,000 purchases. These selection criteria ensure that we can estimate product fixed effects precisely, which helps capture time-invariant differences across products, and that we have sufficient variation to estimate substitution patterns.
For four electronics categories, we collect standard attributes from product detail pages to compare our approach with attribute-based mixed logit: 18 attributes for “Tablets,” 13 for “Monitors,” 20 for “Memory Cards,” and 14 for “Headphones” (see Appendix F for a full list).
@@ 4.2 Estimation Results
We first compare our approach to benchmark models in the four electronics categories for which we collected attribute data. Since we have a relatively large number of attributes in these categories, we reduce their dimensionality via PCA and apply Algorithm 1 for model selection. [^18] We normalize each attribute to have mean zero and variance one before applying PCA to prevent attributes with larger variance or units from dominating the principal components. As in Section 3, we let the candidate set include price and the first $P=6$ principal components.
Table 3 shows the results. In all four categories, unstructured data meaningfully improves model fit relative to both plain logit and attribute-based mixed logit, suggesting that our approach captures information about substitution patterns beyond that contained in standard attributes. This finding is particularly noteworthy given that one might expect technical specifications of electronics products—captured by our extensive list of attributes—to be highly relevant for consumer choices.
[…]
Category: Tablets
Mixed Logit with Attributes 4293.2 -54.9
Mixed Logit with Unstructured Data 4236.6 -111.5
Category: Monitors
Mixed Logit with Attributes 1376.4 -12.3
Mixed Logit with Unstructured Data 1366.4 -22.4
Category: Memory Cards
Mixed Logit with Attributes 2709.3 -72.6
Mixed Logit with Unstructured Data 2685.1 -96.8
Category: Headphones
Mixed Logit with Attributes 9882.0 -4.5
Mixed Logit with Unstructured Data 9875.8 -10.7
Table 3: AICAIC comparison of our approach and attribute-based mixed logit. The table reports the AICAIC of the specification selected by Algorithm 1 within each class of demand models.
These performance differences translate into different substitution patterns. For example, consider diversion ratios for the category of tablets (see Appendix Tables A2-A4). Relative to the plain logit, the selected Description ST model produces more intuitive substitution patterns: kids’ tablets are close substitutes (Fire Kids and Fire HD8 Kids), iPads are close substitutes (iPad 9.7 and iPad 10.2), and so on. The predicted diversions in this model also differ markedly from those in the mixed logit with attributes. However, in the absence of second choice data, we cannot directly test how well different models predict diversions counterfactually, as we do in Section 3. We therefore rely on the fit comparisons reported in Table 3, which favor the Description ST model.
@@
A well-known limitation of the plain logit is that diversion ratios depend only on market shares rather than attribute similarity, leading to overly flat diversions (Conlon et al., 2023). Figure 6 plots the estimated diversion ratio to the closest substitute, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$. The plain logit yields relatively small diversions, about 22% on average, while our approach increases this average to 47% and, in some categories, to as high as 60-80%. These results indicate that our approach produces substantially more variable diversion ratios, suggesting it better identifies close substitutes.